OpenStudy (anonymous):

a body moving with a velocity of 6m/suniformly accelerates at 1.5m/s*s. What will be its velocity after 5 seconds and how far will it move during this time?

random231 (random231):

use v=u+at and s=ut+(at^2)/2

OpenStudy (anonymous):

what is 1.5m/s*s here ? Is it u (Initial velocity) or a (acceleration)

OpenStudy (anonymous):

i got the v=u +at part

random231 (random231):

acceleration

OpenStudy (anonymous):

ok....tysm

OpenStudy (anonymous):

wait a sec, then u becomes - 1.5 m/s

OpenStudy (anonymous):

can it be negative?

random231 (random231):

how can it be negative? its not coming negative check the calculations please!

OpenStudy (anonymous):

STILL NEGATIVE

OpenStudy (anonymous):

Can u show me how u did it?

random231 (random231):



OpenStudy (anonymous):

ohhhhhhh ,thank you very much i thought that 6 was v

random231 (random231):

oh no! v is final velocity and u is the initial velocity